OpenStudy (anonymous):

C = 5/9 (F − 32) Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 8 degrees Celsius is.

OpenStudy (anonymous):

make f as subject 9/5c=f-32 f=?......

OpenStudy (nali):

still need help?

OpenStudy (anonymous):

9/5c=f-32 f=9c/5+32